# Self-supervised prior learning improves structured illumination microscopy resolution

### v1

@@ Introduction
Achieving resolution beyond this twofold enhancement presents a classic ill-posed inverse problem [Tian2025CompSuperRes]. Because the raw data are band-limited by the system’s OTF, there are infinitely many high-resolution structures that could produce the same observed low-resolution image. To determine a unique and physically plausible solution, one must introduce additional constraints, or prior knowledge, about the sample’s structure [Farsiu2004SR, Bouman1993GGMRF, Dey2006RLTV]. Advanced algorithms have incorporated handcrafted priors, such as sparsity and continuity, to improve resolution [Zhao2022SparseSIM], but these priors can be overly simplistic and sensitive to parameter tuning, limiting their generalizability in practice [Chen2023SIMReview].
@@
The deep learning revolution inspired a shift toward data-driven, learnable priors that utilize deep neural networks trained with labeled data for SIM reconstruction [Belthangady2019DLReview]. Notable supervised approaches include U-Net-SIM [Jin2020UNetSIM], DFCAN [Qiao2021DFCAN], and related variants that learn a direct mapping from diffraction-limited or conventionally reconstructed SIM images to super-resolved outputs. However, the practical application of such supervised methods is highly limited by their poor generalization ability when applied to new, unseen testing samples, particularly when the training and test distributions differ in fluorophore density, optical properties, or imaging conditions [Belthangady2019DLReview]. These out-of-distribution generalization issues lead to hallucinations and artifacts, thereby impeding robust generalization in microscopic imaging [Belthangady2019DLReview].
@@
Here we introduce SIMFormer, a Transformer-based [Vaswani2017Attention, Dosovitskiy2021ViT] self-supervised learning framework that resolves the ill-posed challenge of SIM reconstruction by learning a powerful, data-driven prior directly from raw microscopy data. Using a masked autoencoder (MAE) training strategy [He2022MAE, Feichtenhofer2022STMAE], SIMFormer is compelled to learn the intrinsic statistical patterns of biological structures. This learned prior acts as a sophisticated, adaptive regularizer, guiding the reconstruction to solutions that are not only consistent with the measured low-frequency data but also conform to the learned characteristics of plausible biological structures.
@@
Such a process can be viewed as reliable spectral extrapolation, in which the network leverages the available low-frequency information and the strong data-driven prior to infer the most probable high-frequency information that was not directly measured. Consequently, SIMFormer achieves a resolution of approximately 45 nm—nearly a fivefold improvement over the diffraction limit. The framework is fully blind, simultaneously estimating the emitters, illumination patterns, background, and point spread function (PSF) from the raw data alone. To further enhance noise robustness under low-photon-count conditions, we introduce SIMFormer+, a refined model developed through self-distillation [Zhang2019SelfDistill, Hinton2015Distilling, Zhang2022SelfDistillationSurvey, Pham2022RevisitingSelfDistill]. The effectiveness of SIMFormer is thoroughly validated through extensive evaluations on both synthetic and real-world SIM datasets, including the BioSR benchmark [Qiao2021DFCAN], which comprises diverse biological structures such as clathrin-coated pits (CCPs), endoplasmic reticula (ER), microtubules (MTs), and F-actin filaments. Together, these results highlight the broad applicability and practical utility of SIMFormer for super-resolution fluorescence imaging.
@@ Results
Conventional maximum a posteriori (MAP) image reconstruction methods address ill-posedness by incorporating explicit, handcrafted priors such as sparsity and smoothness to regularize the solution [Farsiu2004SR, Bouman1993GGMRF, Dey2006RLTV]. In contrast, SIMFormer utilizes learnable priors that emerge from the data itself through the use of MAE [He2022MAE, Zhang2022MaskMatters]. By reconstructing images from aggressively masked inputs, often discarding around 75% of patches during training, MAE is compelled to capture only the most predictive feature components from the visible patches. This masking-driven process confines the representation to a task-driven, data-adaptive, low-dimensional subspace, effectively creating an implicit prior learned directly from the training data without requiring any manual assumptions on emitters [Zhang2022MaskMatters]. One can formulate the MAE objective in MAP terms as minimizing a reconstruction error with a learned regularizer, where the reconstruction network operates on the target image and the regularizer encodes the learned prior on the latent code.
In SIMFormer, this low-dimensional subspace acts as a powerful prior that constrains the solution to biologically plausible structures. Because the encoder must consistently reconstruct masked regions from the context of unmasked patches, it learns to capture structural regularities such as filament continuity, tubular topology, and network connectivity. These features are precisely the kinds of information needed to disambiguate multiple high-frequency solutions compatible with the same band-limited data. As a result, SIMFormer performs reliable spectral extrapolation while avoiding hallucinations that conflict with the statistics of its training data.
@@
SIMFormer’s approach to generating and utilizing priors fundamentally distinguishes it from other SIM reconstruction methods. Traditional algorithms such as Sparse-SIM [Zhao2022SparseSIM] and Blind-SIM [Mudry2012SpeckleSIM] depend on iterative optimization guided by explicit, handcrafted priors and parametric assumptions about the illumination pattern. These non-learning methods are often computationally intensive and sensitive to parameter tuning [Chen2023SIMReview]. Recent deep learning approaches operate on varied principles. Some, such as the Swin Transformer-based denoising method by Shah et al[Shah2024SwinSIMDenoise] and the self-supervised multimodal denoising method by Chen et al[Chen2024SelfSupDenoiseSIM], are post-process denoisers applied to previously reconstructed images; consequently, they cannot recover spatial frequencies lost during initial reconstruction. Others, for instance, the physics-informed network by Burns and Liu [Burns2023PINNSIM], perform end-to-end reconstruction using an untrained network’s implicit prior. These physics-informed methods typically require known illumination patterns and remain bound by the theoretical $2\times$ resolution limit.
SIMFormer, in contrast, is a full, end-to-end blind reconstruction engine that leverages a strong, data-driven prior learned through self-supervision. This learned prior functions as an adaptive regularizer, allowing for reliable spectral extrapolation beyond conventional resolution limits. Unlike handcrafted priors that must be manually tuned for different structures and imaging conditions, the learned prior in SIMFormer emerges directly from raw training data and adapts to complex biological structures. Moreover, by jointly estimating emitters, patterns, background, and PSF, SIMFormer is able to account for experimental imperfections that are hard to model analytically, such as deviations from ideal sinusoidal illumination and subtle aberrations in the PSF.
Network architecture of SIMFormer
SIMFormer introduces a novel architecture for fully blind SIM reconstruction with four main components.
(1) Data processor: Raw SIM images are taken as an input stack consisting of three phases from each of three SIM directions. The input stack is divided into $3\times 16\times 16$ 3D cubes, which are transformed into 768-dimensional vectors by patch embedding.
(2) Encoder: The embedded data are processed by a Transformer encoder derived from ViT [Dosovitskiy2021ViT] (ViT-base) with 12 Transformer layers. To accelerate training and improve dense prediction capabilities, we introduce an adapter module into the encoder [Chen2023ViTAdapter]. The encoder is initialized by CLIP-based pre-training [Radford2021CLIP] to leverage large-scale natural image representations and improve optimization stability.
(3) Decoder: The encoded features are processed by a decoder with multiple Transformer layers. The decoder output is passed through three convolutional decoding heads to separately reconstruct emitters, light patterns, and background, which together synthesize the raw SIM images. For the light pattern, a low-rank coding module is introduced before the convolutional decoder head to progressively compress the rank of the light pattern representation during training, reflecting the prior knowledge that only a few pattern modes are physically present.
(4) PSF branch: The PSF is modeled by feeding a fixed random embedding to a learnable convolutional neural network, following the philosophy of deep image priors [Ulyanov2018DeepImagePrior, Wang2024ImagingMechanism]. Overall, the raw SIM images are synthesized using the emitters, light patterns, background, and PSF according to the physical principles of SIM imaging.
The entire pipeline (Fig. 1) is trained in a self-supervised manner: the network takes raw SIM stacks as input and predicts emitters, patterns, background, and PSF whose forward simulation must match the original raw stacks. This design tightly couples the learned prior with a physically realistic forward model.
@@
Our SIMFormer adopts a multi-stage self-supervised training approach inspired by curriculum learning [Bengio2009Curriculum]. At each stage, the masking rate and the compression ratio of the low-rank coding are adjusted: the masking rate decreases, while the compression ratio gradually increases, following a multi-stage schedule. This design aims to prevent the model from converging to trivial PSF solutions in the early stages and to promote a stable progression toward low-rank coding for the light pattern.
The training objective consists of four parts: (1) a reconstruction loss measuring the difference between synthesized and raw SIM images; (2) a total variation loss enforcing smoothness in the predicted light pattern; (3) a total variation loss for the PSF to encourage smooth, physically plausible responses; and (4) a mean absolute error loss enforcing alignment between the PSF centroid and the image center. A weighted combination of $L_{1}$ and MS-SSIM terms is used for image fidelity, as is common in image restoration networks [Zhao2016LossFunctions]. Details are provided in the Materials and Methods section.
Self-distillation and SIMFormer+
Knowledge distillation [Hinton2015Distilling] trains a student model to replicate the behavior of a teacher model. We adopt a self-distillation approach [Zhang2019SelfDistill, Zhang2022SelfDistillationSurvey, Pham2022RevisitingSelfDistill], where the student model shares the same architecture as the teacher. Using the teacher’s predictions as soft labels, the student model refines and enhances the original outputs, resulting in improved accuracy and stronger noise robustness. This refined model is referred to as SIMFormer+.
Specifically, SIMFormer is first trained on high-SNR SIM images, producing high-resolution predictions. These predictions are then used as pseudo ground truth to supervise SIMFormer+ at lower photon counts and higher noise levels. By matching the teacher’s soft predictions rather than hard labels, the student model learns to preserve fine structures while being explicitly optimized for robustness under challenging noise conditions. As a result, SIMFormer+ inherits the teacher’s high-resolution capabilities while substantially improving robustness to photon starvation and detector noise.
@@
We demonstrate SIMFormer’s capabilities using both synthetic and real-world data. For synthetic validation, several datasets with varying fluorophore structures, light pattern spatial periods, noise levels, line sparsity, tube widths, and ring radii are generated. In these datasets, the fluorescence wavelength is set to 488 nm, the objective numerical aperture (NA) is 1.3, and the pixel size is 62.6 nm. The PSF is simulated using a vectorial Debye integral model, corresponding to a diffraction limit of 229 nm. The light pattern is modeled as a cosine function oriented in three directions with three phases per direction. Using the emitters, light patterns, and PSF, we synthesize raw SIM images and reconstruct super-resolution images using a conventional algorithm (Supplementary Notes 2 and 3). For each parameter configuration, we generate 100 training samples and 100 test samples.
For real-world validation, we adopt the BioSR dataset [Qiao2021DFCAN], which contains raw SIM images of various biological structures, including CCPs, ER, MTs, and F-actin, along with the corresponding super-resolution images obtained through conventional SIM reconstruction. This dataset covers a wide range of structural densities and morphologies, providing a stringent test for reconstruction algorithms.
@@
We evaluate SIMFormer using both synthetic and real experimental data and compare it with conventional SIM reconstruction, Sparse-SIM [Zhao2022SparseSIM], and Blind-SIM [Mudry2012SpeckleSIM]. Sparse-SIM is currently one of the state-of-the-art SIM reconstruction algorithms with high resolution [Zhao2022SparseSIM], while Blind-SIM is a widely used blind reconstruction algorithm [Mudry2012SpeckleSIM]. Using the labeled synthetic dataset, we examine the superiority of SIMFormer over Sparse-SIM and Blind-SIM under two challenging conditions: (1) SIMFormer is trained and predicts in a fully blind manner, whereas Sparse-SIM has access to experimental parameters and Blind-SIM uses the true PSF; and (2) the performance of Sparse-SIM and Blind-SIM is fine-tuned using global optimization of their hyperparameters (Supplementary Note 4), whereas SIMFormer is not fully optimized due to computational resource limitations. In real-world experiments, Blind-SIM uses the same parameters as in the synthetic experiments, while Sparse-SIM employs new parameters (Supplementary Table 2).
@@
The synthetic dataset provides actual emitters as ground truth, allowing the use of normalized root-mean-square error (NRMSE) and multiscale structural similarity index (MS-SSIM) to evaluate reconstructed images [Zhao2016LossFunctions]. NRMSE quantifies overall intensity differences between predicted and actual images, while MS-SSIM assesses structural similarity across multiple scales. NRMSE with added Gaussian blur is used to measure spatial distance error of emitters (Supplementary Note 5). Since the resolution achieved by Sparse-SIM and SIMFormer surpasses conventional SIM reconstruction in the BioSR dataset (where no ground truth is available), we focus on discriminability of regions of interest and compare results with prior knowledge from STORM imaging of ERs [Wang2022ERTubules]. Additionally, decorrelation analysis [Descloux2019Decorrelation] is used to measure resolution (Supplementary Note 6). For noise robustness evaluation, we use fairSIM [Muller2016fairSIM] to generate conventional SIM reconstructions under high noise, as these results are not available in BioSR.
SIMFormer on synthetic data
We first evaluate SIMFormer on synthetic filament structures, where fluorophores are simulated using randomly generated curves based on Bessel functions (Supplementary Note 2). SIMFormer and SIMFormer+ substantially improve NRMSE and MS-SSIM compared with widefield imaging, conventional SIM, Blind-SIM, and Sparse-SIM, and achieve significantly improved resolution according to decorrelation analysis [Descloux2019Decorrelation]. Zoomed-in regions show that: (1) conventional SIM fails to resolve fine buckle structures; (2) Blind-SIM introduces significant artifacts; (3) Sparse-SIM shows fewer artifacts but still cannot distinctly resolve the structures; (4) SIMFormer produces plausible reconstructions with fewer artifacts; and (5) SIMFormer+ yields the clearest buckle structures owing to self-distillation improving robustness.
We then evaluate synthetic microtubule structures. SIMFormer and SIMFormer+ again outperform existing methods in both NRMSE and MS-SSIM, and achieve higher resolution than Blind-SIM and Sparse-SIM. In zoomed-in regions, microtubule structures narrower than the diffraction limit are resolved only by SIMFormer and SIMFormer+. Extended Data Fig. 2 shows that Sparse-SIM and Blind-SIM can resolve only wider microtubules, with Sparse-SIM performing relatively better than Blind-SIM.
@@
Finally, we simulate ring-shaped fluorophores with varying radii. SIMFormer and SIMFormer+ again significantly improve NRMSE and MS-SSIM, and decorrelation analysis indicates that SIMFormer+ achieves about 2.8 times the resolution of conventional SIM and 1.5 times that of Blind-SIM and Sparse-SIM. SIMFormer and SIMFormer+ resolve ring structures of about 75 nm diameter, which conventional SIM, Sparse-SIM, and Blind-SIM do not. Extended Data Fig. 3 shows that Sparse-SIM resolves rings 125 nm or larger, while Blind-SIM resolves rings 165 nm or larger. Overall, considering discernibility, accuracy, and resolution, SIMFormer+ achieves the best performance across all synthetic tasks.
SIMFormer on the BioSR dataset
@@
We next use the BioSR dataset [Qiao2021DFCAN] to evaluate SIMFormer in real-world conditions. The dataset contains CCPs, ER, MTs, and F-actin filaments, covering structures from sparse to dense. With fluorescence wavelength 488 nm and NA 1.3, the diffraction limit is 229 nm. Raw SIM images are processed using Blind-SIM, Sparse-SIM, and SIMFormer, alongside conventional SIM reconstructions provided in BioSR. Self-distillation on SIMFormer results yields SIMFormer+.
Fig. 3 shows a side-by-side comparison of methods on BioSR. CCPs are ring-like assemblies of clathrin, typically 50–150 nm in diameter and essential for endocytosis [McMahon2011Clathrin]. SIMFormer is the only method that detects sub-100 nm micro-rings, while all methods detect larger rings; SIMFormer+ matches SIMFormer and further suppresses residual artifacts.

### v2

@@ Introduction
Achieving resolution beyond this twofold enhancement presents a classic ill-posed inverse problem [Tian2025CompSuperRes]. Because the raw data are band-limited by the system’s OTF, there are infinitely many high-resolution structures that could produce the same observed low-resolution image. To determine a unique and physically plausible solution, one must introduce additional constraints, or prior knowledge, about the sample’s structure [Tian2025CompSuperRes]. Advanced algorithms have incorporated handcrafted priors, such as sparsity and continuity, to improve resolution [Zhao2022SparseSIM], but these priors can be overly simplistic and sensitive to parameter tuning, limiting their generalizability in practice [Chen2023SIMReview].
@@
The deep learning revolution inspired a shift toward data-driven learnable priors, utilizing deep neural networks trained with labeled data for SIM reconstruction. Notable supervised approaches include U-Net-SIM [Jin2020UNetSIM], DFCAN [Qiao2021DFCAN], and DFGAN [Qiao2021DFCAN], employ supervised learning approaches. These methods rely on high-resolution images—either simulated or experimentally measured with other microscopy techniques—to serve as ground truth for training. However, the practical application of such supervised methods is highly limited by their poor generalization ability when applied to new, unseen testing samples [Belthangady2019DLReview, Weigert2018CARE]. This lack of adaptability stems from out-of-distribution generalization issues, where discrepancies between the labeled training data and new experimental data undermine the reliability of the results.
@@
Here we introduce SIMFormer, a Transformer-based [Vaswani2017Attention, Dosovitskiy2021ViT] self-supervised learning framework that resolves the ill-posed challenge of SIM reconstruction by learning a powerful, data-driven prior directly from raw microscopy data. Using a masked autoencoder (MAE) [He2022MAE, Feichtenhofer2022STMAE] training strategy, SIMFormer is compelled to learn the intrinsic statistical patterns of biological structures. This learned prior acts as a sophisticated, adaptive regularizer, guiding the reconstruction to solutions that are not only consistent with the measured low-frequency data but also conform to the learned characteristics of plausible biological structures.
@@
Such a process can be viewed as reliable spectral extrapolation, in which the network leverages the available low-frequency information and the strong data-driven prior to infer the most probable high-frequency information that was not directly measured. Consequently, SIMFormer achieves a resolution of approximately 45 nm—nearly a fivefold improvement over the diffraction limit. The framework is fully blind, simultaneously estimating the emitters, illumination patterns, background, and point spread function (PSF) from the raw data alone. To further enhance noise robustness under low-photon-count conditions, we introduce SIMFormer+, a refined model developed through self-distillation [Zhang2019SelfDistill]. The effectiveness of SIMFormer is thoroughly validated through extensive evaluations on both synthetic and real-world SIM datasets, including the BioSR benchmark [Qiao2021DFCAN], which comprises diverse biological structures such as clathrin-coated pits (CCPs), endoplasmic reticula (ER), microtubules (MTs), and F-actin filaments. Together, these results highlight the broad applicability and practical utility of SIMFormer for super-resolution fluorescence imaging.
@@ Results
Conventional maximum a posteriori (MAP) image reconstruction methods address ill-posedness by incorporating explicit, handcrafted priors such as sparsity and smoothness to regularize the solution [Farsiu2004SR, Bouman1993GGMRF, Dey2006RLTV]. In contrast, SIMFormer utilizes learnable priors that emerge from the data itself through the use of MAE [He2022MAE]. By reconstructing images from aggressively masked inputs, often discarding around 75% of patches during training, MAE is compelled to capture only the most predictive feature components from the visible patches [Zhang2022MaskMatters]. This masking-driven process confines the representation to a task-driven, data-adaptive, low-dimensional subspace, effectively creating an implicit prior learned directly from the training data without requiring any manual assumptions on emitters [Zhang2022MaskMatters]. One can formulate the MAE objective in MAP terms as minimizing a reconstruction error with a learned regularizer: $\min_{z}\|D(z)-x\|_{2}^{2}+\lambda R(z)$, where $D(z)$ reconstructs the target image $x$ and $R(z)$ encodes the learned prior on the latent code.
@@
SIMFormer’s approach to generating and utilizing priors fundamentally distinguishes it from other SIM reconstruction methods. Traditional algorithms such as Sparse-SIM [Zhao2022SparseSIM] and Blind-SIM [Mudry2012SpeckleSIM] depend on iterative optimization guided by explicit, handcrafted priors and parametric assumptions about the illumination pattern. These non-learning methods are often computationally intensive and sensitive to parameter tuning [Chen2023SIMReview]. Recent deep learning approaches operate on varied principles. Some, such as the Swin Transformer-based denoising method by Shah et al[Shah2024SwinSIMDenoise] and the self-supervised multimodal denoising method by Chen et al[Chen2024SelfSupDenoiseSIM], are post-process denoisers applied to previously reconstructed images; consequently, they cannot recover spatial frequencies lost during initial reconstruction. Others, for instance, the physics-informed network by Burns and Liu [Burns2023PINNSIM], perform full reconstruction but utilize an untrained network’s implicit prior. These methods require known illumination patterns and are bound by the theoretical resolution limit. SIMFormer, in contrast, is a full, end-to-end blind reconstruction engine that leverages a strong, data-driven prior learned through self-supervision. This learned prior functions as an adaptive regularizer, allowing for reliable spectral extrapolation beyond conventional resolution limits, a capability not provided by the principles of other methods.
The network architecture of SIMFormer
SIMFormer introduces a novel architecture for fully blind SIM reconstruction which contains four components. (1) Data processor. It takes raw SIM images as an input stack consisting of the three phases from each of the SIM image’s three directions. The input stack is divided into $3\times 16\times 16$ 3D cubes, and then they are transformed into 768-dimensional vectors by patch embedding. (2) Encoder. The embedded data is further processed by a transformer encoder derived from ViT [Dosovitskiy2021ViT] (ViT-base) with 12 transformer layers. Meanwhile, an adapter is introduced into the encoder to accelerate training and effectively improve dense prediction capabilities [Chen2023ViTAdapter]. (3) Decoder. The encoded features are processed by a decoder consisting of 12 transformer layers, the output of which is passed through three convolutional decoding heads to separately output emitters, light patterns, and background to reconstruct the raw SIM images. Particularly, for the light pattern, a low-rank coding module is introduced before the convolutional decoder head to progressively compress the rank of the light pattern representation during training. (4) PSF. It is modeled by feeding a fixed random embedding to a learnable convolutional neural network. Overall, the raw SIM images are synthesized using the emitters, light patterns, background, and PSF following the principles of SIM imaging. See Extended Data Fig. 1 for details.
@@
Our SIMFormer adopts a multi-stage self-supervised training approach inspired by curriculum learning [Bengio2009Curriculum]. At each stage, the masking rate and the compression ratio of the low-rank coding will be adjusted. Specifically, the masking rate decreases, while the compression ratio gradually increases, following a multi-stage schedule. This design aims to prevent the model from converging to trivial PSF solutions in the early stages and promote a stable progression toward low-rank coding for the light pattern.
Our training objective involves four parts as follows: (1) a reconstruction loss function to measure the difference between the synthesized and raw SIM images; (2) a total variation loss function to enforce smoothness in the predicted light pattern; and (3) a total variation loss function for PSF and light pattern smoothness, along with (4) a mean absolute error loss to ensure alignment between the PSF centroid and the image center. See Methods for further details.
Self Distillation
Knowledge distillation [Hinton2015Distilling] is a machine learning technique where a student model is trained to replicate the behavior of a teacher model. In this study, we adopt a self-distillation [Zhang2019SelfDistill, Zhang2022SelfDistillationSurvey, Pham2022RevisitingSelfDistill] approach, where the student model shares the same architecture as the teacher. By using the teacher’s predictions as soft labels, the student model learns to refine and enhance the original outputs, resulting in improved accuracy and stronger noise robustness. This refined model is referred to as SIMFormer+.
@@
We demonstrate SIMFormer’s capabilities using both synthetic and real-world data. Several datasets with varying fluorophore structures, light pattern spatial periods, noise levels, line sparsity, tube widths, and ring radii are generated for synthetic validation. In these datasets, the fluorescence wavelength is set to 488 nm, the objective’s numerical aperture (NA) is 1.3, and the pixel size is 62.6 nm. The PSF is simulated using the Debye integral method [Zhang2007GaussianPSF], corresponding to a diffraction limit of 229 nm. The light pattern is modeled as a cosine function, oriented in three directions with three phases for each direction. Using the emitters, light patterns, and the PSF, we synthesized raw SIM images and subsequently reconstructed super-resolution images using conventional algorithm (Supplementary Note 2 and 3). For each parameter configuration, we generated 100 samples for the training set and another 100 samples for the test set. As for real-world validation, the BioSR dataset is adopted, which contains raw SIM images of various biological structures, including CCPs (54 cells), ER (68 cells), MTs (55 cells), and F-actin (51 cells), along with the corresponding super-resolution images obtained through the conventional SIM reconstruction algorithm.
@@
We use both synthetic and real experimental data to evaluate the performance of SIMFormer. Meanwhile, it is compared with representative existing approaches including conventional SIM reconstruction, Sparse-SIM [Zhao2022SparseSIM], and Blind-SIM [Mudry2012SpeckleSIM]. Sparse-SIM is currently the state-of-the-art SIM reconstruction algorithm with the highest resolution [Chen2023SIMReview], while Blind-SIM is a widely used blind reconstruction algorithm [Lee2021MetamaterialNanoscopy]. Using the labeled synthetic dataset, we examine the superiority of SIMFormer over Sparse-SIM and Blind-SIM under two challenging conditions: (1) SIMFormer is trained and makes predictions in a fully blind manner, whereas Sparse-SIM has access to experimental parameters, and Blind-SIM uses the true PSF; (2) The performance of Sparse-SIM and Blind-SIM is fine-tuned using genetic algorithms to achieve optimal results (Supplementary Note 4). However, SIMFormer has not been optimized due to computational resource limitations. In real-world experiments, Blind-SIM used the same parameters as in the synthetic data experiments, while Sparse-SIM employed new parameters to achieve better performance (Supplementary Note Table 2).
@@
The synthetic dataset provides actual emitters as ground truth, allowing us to use the normalized root-mean-square error (NRMSE) and the multiscale structural similarity index (MS-SSIM) to evaluate the reconstructed images. NRMSE quantifies the overall difference in intensity between the predicted and actual images, providing a measure of the error magnitude. In contrast, MS-SSIM assesses the structural similarity between the images at multiple scales, focusing on perceived visual quality related to structural information. Note that we use NRMSE with added Gaussian blur to measure the spatial distance error of the emitters (Supplementary Note 5). Since the resolution achieved by Sparse-SIM and SIMFormer surpasses that of conventional SIM reconstruction in the BioSR dataset (so we don’t have ground truth), we focus on the discriminability of regions of interest and compare the results with prior knowledge from STORM imaging of ERs [Wang2022ERTubules]. Additionally, we apply decorrelation analysis [Descloux2019Decorrelation] to measure resolution (Supplementary Note 6). For the noise robustness evaluation, we use fairSIM [Muller2016fairSIM] to generate conventional SIM reconstruction results under high noise levels, as these results are not available in BioSR.
SIMFormer demonstrates highly accurate super-resolution imaging on synthetic data
We first evaluate SIMFormer on synthetic filament structures, where fluorophores are simulated using randomly generated curves based on Bessel functions (Supplementary Note 2). As shown in Fig. 2d, the NRMSE results are as follows: Widefield 1.15, conventional SIM 0.87, Blind-SIM 0.67, Sparse-SIM 0.59, SIMFormer 0.32, and SIMFormer+ 0.30. Notably, SIMFormer+ achieves a 49.5% improvement in NRMSE over Sparse-SIM and a 55.7% improvement over Blind-SIM. For MS-SSIM, the values are: Widefield 0.64, conventional SIM 0.63, Blind-SIM 0.75, Sparse-SIM 0.85, SIMFormer 0.90, and SIMFormer+ 0.92. Here, SIMFormer+ demonstrates an 8.2% improvement over Sparse-SIM and a 22.3% improvement over Blind-SIM. As shown in Fig. 2g, the resolution values from decorrelation analysis indicate: Widefield 219.4 nm, conventional SIM 122.0 nm, Blind-SIM 67.3 nm, Sparse-SIM 69.2 nm, SIMFormer 48.5 nm, and SIMFormer+ 49.3 nm. SIMFormer+ achieves approximately 2.5 times the resolution of conventional SIM and 1.4 times better resolution than both Blind-SIM and Sparse-SIM.
A closer examination of the zoomed-in regions, particularly the 46 nm buckle structure, highlights several important observations that support the quantitative results: (1) conventional SIM reconstruction fails to resolve the buckle structure, (2) Blind-SIM reconstruction implies a rough structure but introduces significant artifacts, (3) Sparse-SIM shows fewer artifacts, yet still cannot distinctly resolve the structure, (4) SIMFormer produce a rough buckle structure with fewer reconstruction artifacts, attributed to the learned priors, and (5) SIMFormer+ produce a clear buckle structure compared to its vanilla counterpart, thanks to our self-distillation designs which brings more reconstruction robustness. Moreover, Supplementary Fig. 1 presents the reconstruction results for samples varying from sparse to dense, where SIMFormer and SIMFormer+ show a significant performance advantage, particularly in dense samples.
Next, we evaluate SIMFormer on synthetic microtubule structures. As shown in Fig. 2e, the NRMSE results are: Widefield 1.16, conventional SIM 0.83, Blind-SIM 0.70, Sparse-SIM 0.61, SIMFormer 0.34, and SIMFormer+ 0.33. SIMFormer+ demonstrates a 47.0% improvement over Sparse-SIM and a 53.2% improvement over Blind-SIM. For MS-SSIM, the values are: Widefield 0.56, conventional SIM 0.59, Blind-SIM 0.74, Sparse-SIM 0.82, SIMFormer 0.86, and SIMFormer+ 0.91, where SIMFormer+ achieves a 10.5% improvement over Sparse-SIM and a 22.2% improvement over Blind-SIM. As shown in Fig. 2h, resolution values from decorrelation analysis indicate: Widefield 238.3 nm, conventional SIM 127.8 nm, Blind-SIM 66.3 nm, Sparse-SIM 73.3 nm, SIMFormer 55.3 nm, and SIMFormer+ 51.7 nm. SIMFormer+ achieves approximately 2.5 times the resolution of conventional SIM, 1.3 times better resolution than Blind-SIM, and 1.4 times better resolution than Sparse-SIM.
In Fig. 2b, the zoomed-in regions highlight microtubule structures that are 65 nm wide. These structures are not clearly resolved by any method except SIMFormer and SIMFormer+, as they exceed the recognition limits of blind reconstruction and hand-crafted priors, leading to ambiguity. Extended Data Fig. 2 shows that while both Sparse-SIM and Blind-SIM can resolve wider microtubule structures (approximately 126 nm), Sparse-SIM performs relatively better than Blind-SIM.
@@
Finally, we simulate ring-shaped fluorophores with varying radii to validate SIMFormer on ring structures. As shown in Fig. 2f, the NRMSE results are: Widefield 1.19, conventional SIM 0.85, Blind-SIM 0.65, Sparse-SIM 0.65, SIMFormer 0.32, and SIMFormer+ 0.29. SIMFormer+ improves NRMSE by 55.9% compared to Sparse-SIM and Blind-SIM. For MS-SSIM, the values are: Widefield 0.68, conventional SIM 0.70, Blind-SIM 0.85, Sparse-SIM 0.88, SIMFormer 0.91, and SIMFormer+ 0.95, showing a 7.9% improvement over Sparse-SIM and 10.8% over Blind-SIM. Furthermore, resolution values from decorrelation analysis in Fig. 2i indicate: Widefield 238.7 nm, conventional SIM 135.3 nm, Blind-SIM 71.9 nm, Sparse-SIM 72.3 nm, SIMFormer 48.7 nm, and SIMFormer+ 47.9 nm. SIMFormer+ achieves 2.8 times the resolution of conventional SIM and 1.5 times better resolution than both Blind-SIM and Sparse-SIM. In Fig. 2c, SIMFormer and SIMFormer+ successfully resolve ring structures about 75 nm in diameter, which conventional SIM, Sparse-SIM, and Blind-SIM do not. Extended Data Fig. 3 shows Sparse-SIM resolving rings 125 nm or larger, while Blind-SIM resolves rings 165 nm or larger.
Overall, considering discernibility, accuracy, and resolution, the ranking of methods is SIMFormer+, SIMFormer, Sparse-SIM, Blind-SIM, and conventional SIM. Quantitative and visual metrics consistently support this ranking assessment across all benchmarks. These performance differences highlight a fundamental distinction in how each method addresses the ill-posed nature of SIM reconstruction. Methods like Sparse-SIM impose hand-crafted priors, such as sparsity, which can be overly rigid for complex structures. In contrast, SIMFormer utilizes a flexible and powerful prior learned directly from data. This self-supervised approach forces the model to understand the underlying statistical patterns of microscopy images, thereby acting as an adaptive regularizer. Consequently, when faced with ambiguous regions, such as the dense 46 nm buckle, SIMFormer resolves the structure more reliably, avoiding the tendency of methods like Sparse-SIM to interpret such areas as fine lines.
SIMFormer achieves 45 nm resolution on the BioSR dataset
@@
Here, we use the BioSR dataset to evaluate SIMFormer in real-world conditions. The BioSR dataset consists of four structure types—CCPs, ER, MTs, and F-actin filaments—ranging from sparse to dense. With a fluorescence wavelength of 488 nm and an NA of 1.3, the diffraction limit is 229 nm. We process the raw SIM images using Blind-SIM, Sparse-SIM, and SIMFormer, alongside conventional SIM reconstructions provided in the BioSR dataset. Then, self-distillation is performed on the SIMFormer results to obtain SIMFormer+. As shown in Fig. 3, these results are compared side-by-side with the wide-field imaging results.
CCPs are specialized regions of the cell membrane responsible for endocytosis, where cells internalize essential molecules such as nutrients and receptors. Structurally, CCPs are ring-like assemblies of the protein clathrin, typically measuring 50 to 150 nm in diameter, and their size is crucial for vesicle formation and cargo selection [McMahon2011Clathrin]. Fig. 3a presents the comparison results for CCPs. In previous synthetic data experiments, SIMFormer demonstrated its capability to resolve 75 nm ring structures. In this case, only SIMFormer can detect the 78 nm micro-ring, while all methods identify the 125 nm micro-ring. SIMFormer+ achieves consistent results with SIMFormer. Additional results are shown in Supplementary Fig. 2.